Given the individual consumption function C = $1000 + 0.95YD,

Instructions: Enter your responses as a whole number.

a. How much is saved if YD is $60,000?

$

b. How much does consumption increase if YD increases by $1,000?

$
Answer #1

C = $1000 + 0.95YD

a.

C = 1000 + (0.95*60,000)

C = 58,000

Savings = YD - Consumption

Savings = 60,000 - 58,000

Savings = $2,000

b.

MPC = 0.95

Multiplier = 1 / 1 - MPC

= 1 / 1 - 0.95

= 20

Multiplier = Change in Income / change in Consumption

Change in consumption = Change in Income / Multiplier

Change in consumption = 1000 / 20

Change in consumption = 50

So, consumption increases by $50

**if you liked the answer, then please upvote. Would be motivating for me. Thanks.